how do you solve a question like: brian is 26 years older than his daughter kim. the sum of their ages is twice the difference of their ages. find their ages.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Use b & k for the ages to avoid typing.
---------------
Write down what you're told:
b = k + 26 (he's 26 years older)
------
The sum of their ages is b + k
The difference is b - k
------
Then, (b+k) is 2 times (b-k)
b+k = 2(b-k)
Sub (k+26) for b, solve for k
etc

Answer by stanbon(75887)

brian is 26 years older than his daughter kim. the sum of their ages is twice the difference of their ages. find their ages.
-----
Equations:
b = k+26
b + k = 2(b-k)
-------------------------
Substitute for "b" and solve for "k":
k+26 +k = 2(k+26-k)
---
2k + 26 = 52
2k = 26
k = 13 (Kim's age)
-----------------------
Since b = k+26, b = 39 yrs (brian's age)
===========================================
Cheers,
Stan H.